OpenStudy (anonymous):

what is 40/20 in a decimal

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Simplify as a fraction; 40/20 = 4/2 = 2/1 Then, 2 divided by 1 = 2 (Or 2.0 as a decimal) Hope I helped! ;D